OpenStudy (anonymous):

the average of 4 numbers is 6 if one of the numbers is 5 what is the sum of the other 3 members? A.24 B.19 C.15 D.1

OpenStudy (anonymous):

5+(x+y+z)/4=6 x+y+z=24-5 x+y+z=19